OpenStudy (anonymous):

Suppose you wish to conduct a study on the average length of a hospital stay. You would like to estimate the average length to within 0.5 days of the true mean at 90% confidence. Assuming σ=3.7, how many patients should you randomly select? What if you wanted 95% confidence?

OpenStudy (amistre64):

\[E=z\sqrt{\frac{pq}{n}}\] does that seem right?

OpenStudy (amistre64):

solving for n we get \[n=\frac{z^2pq}{E^2}\]

OpenStudy (amistre64):

or maybe this is one of those z = x-mean/(modified sd)

OpenStudy (amistre64):

\[z=\frac{x-\bar x}{sd/\sqrt{n}}\]

OpenStudy (anonymous):

I think that the first equation you gave me might be right but I'm confused because it's asking me for "n" when it hasn't given me the mean.

OpenStudy (amistre64):

is this a 2 part question perhaps? or do we keep the mean generic?

OpenStudy (anonymous):

Normally with this type of question I would do Z=mean+-Z*(SD/sqrt of n) and then solve for n

OpenStudy (amistre64):

z is just the zscore of the given CI ... \[z=\frac{x-\mu}{\sigma/\sqrt{n}}\] \[\frac{z\sigma}{\sqrt{n}}=x-\mu\] \[\frac{z\sigma}{x-\mu}=\sqrt{n}\] \[\left(\frac{z\sigma}{x-\mu}\right)^2=n\]

OpenStudy (amistre64):

assume a mean of 0 ...

OpenStudy (anonymous):

Okay, thank you for your help!

OpenStudy (amistre64):

x=.5 , mu = 0 , sigma = 3.7 , z = zscore of 90% and then 95%

OpenStudy (amistre64):

" "True" Mean and Confidence Interval. Probably the most often used descriptive statistic is the mean. The mean is a particularly informative measure of the "central tendency" of the variable if it is reported along with its confidence intervals. As mentioned earlier, usually we are interested in statistics (such as the mean) from our sample only to the extent to which they can infer information about the population. The confidence intervals for the mean give us a range of values around the mean where we expect the "true" (population) mean is located (with a given level of certainty, see also Elementary Concepts). For example, if the mean in your sample is 23, and the lower and upper limits of the p=.05 confidence interval are 19 and 27 respectively, then you can conclude that there is a 95% probability that the population mean is greater than 19 and lower than 27. If you set the p-level to a smaller value, then the interval would become wider thereby increasing the "certainty" of the estimate, and vice versa; as we all know from the weather forecast, the more "vague" the prediction (i.e., wider the confidence interval), the more likely it will materialize. Note that the width of the confidence interval depends on the sample size and on the variation of data values. The larger the sample size, the more reliable its mean. The larger the variation, the less reliable the mean (see also Elementary Concepts). The calculation of confidence intervals is based on the assumption that the variable is normally distributed in the population. The estimate may not be valid if this assumption is not met, unless the sample size is large, say n=100 or more." http://www.statsoft.com/textbook/basic-statistics/#Descriptive statisticsa